Suppose random variables X and Y are related as Y=7X+3. Suppose the random variable X has mean 1, and variance 1. What is the expected value of X+Y ?

Solution provided is 11. How was this obtained?
Answer #1

we have given Y = 7X+3

and E(X) = mean = 1

VAR(X) = 1

E(Y) = E(7X+3) = 7E(X)+3 = 7*1+3 = 10

now E(X+Y) =E(X) +E(Y) = 1+10 = 11